OpenStudy (anonymous):

AP Calculus AB Question:

OpenStudy (anonymous):

Find the following integral using only methods from Calc AB (no integration by parts, etc.) \[\int\limits_{-2}^{2}\frac{ 1 }{ x^2+4 } dx\]

OpenStudy (experimentx):

\[ \frac{d}{dx} \arctan(x) = \frac{1}{1+x^2}\] make suitable substitution.

OpenStudy (anonymous):

actually the 1 should be a 4, how do I do it now?

OpenStudy (experimentx):

take 4 common ... substitute x/2 = y

OpenStudy (anonymous):

I don't think I'm getting it, I'm not too great at these, could you show me a worked out solution if you have time?

OpenStudy (experimentx):